OpenStudy (anonymous):

Find the linearization L(x) of the function at a. 1. F(x)=x^(3/4), a=16

OpenStudy (anonymous):

L(x)=F(a) + F'(a)*(x-a) F(a)=16^(3/4) F'(x)=3(x^(-1/4))/4 F'(a)=3/4 * 16^(-1/4) L(x)=16^(3/4) + (3/4 * 16^(-1/4))(x-16) L(x)=8+(3/8)(x-16)